### v1

# Diameter Reduction Via Flipping Arcs

Abstract
The diameter of a directed graph is a fundamental parameter defined as the maximum distance realized among the pairs of vertices. As graphs of small diameter are of interest in many applications, we study the following problem: for a given directed graph and a positive integer $d$, what is the minimum number of arc flips (also known as arc reversal) required to obtain a graph with diameter at most $d$? It is a generalization of the well-known problem Oriented Diameter, first studied by Chvátal and Thomassen.
Here we investigate variants of the above problem, considering the number of flips and the target diameter as parameters. We prove that most of the related questions of this type are hard. Special cases of graphs are also considered, as planar and cactus graphs, where we give polynomial time algorithms.
Keywords: diameter, diameter reduction, edge flip, arc reversal, edge polytope
1 Introduction
The well-known theorem of Robbins [16] states that an undirected graph admits a strongly connected orientation, if and only if it is 2-edge-connected [^1] An undirected graph is ($k$-edge-connected)$k$-arc-connected if at least $k$ (edges)arcs have to be removed in order to destroy connectivity.. His theorem was inspired by an application in road traffic control, to make traffic flow more efficient when using only one-way streets. However, what makes an orientation good can vary according to the problem description. For example, we might want to minimize the distance from a given vertex to all other vertices or we might seek to minimize the average distance between any pair of vertices. Here, we consider an oriented diameter problem on directed graphs which has turned out to be a useful concept and it now has a rich literature. In this problem our goal is to minimize the maximal distance realized between the vertices (called the diameter of the graph) by changing the direction of as few arcs as possible. One motivation of the problem is the following: Suppose that there are $n$ airports, some ordered pairs are connected by a flight. For a given integer $d$ how many flights do we have to established to make sure that from each airport one can get to another by changing planes at most $d$ times?
Knowing the diameter of a graph is very helpful in many applications for example in analysis of social networks, in parallel computing and even in linguistics. To formalise the problem more precisely, we introduce some notation. Given a graph (directed or undirected) $D=(V,A)$, we denote an edge between $u$, $v\in V$ by $uv$ and an arc from $u$ to $v$ by $(u,v)$. Also, we denote a path $P$ between $u,v\in V$ by $(u,w_{1},w_{2},\dots,w_{\ell},v)$ where $w_{0}=u,w_{1},\dots,w_{\ell},w_{\ell+1}=v$ are the nodes of the path so that $(w_{i},w_{i+1})\in A$ for all $i=0,\dots,\ell$. The length of $P$ is the number of edges (or arcs) and it is denoted by $|P|$.
For any $u,v\in V$, the distance between them is the length (or weight if the edges are weighted) of the shortest path connecting $u$ and $v$; by convention, this distance is $\infty$ if there is no path connecting the vertices. The diameter of the graph $D$, denoted by $\operatorname{diam}(D)$ is the maximum distance realized among its vertices. An arc reversal of $e=(u,v)$ in a directed graph $G=(V,A)$ results on a graph $(V,(A\setminus(u,v))\cup(v,u))$. We also call it a flipping. Let $F\subseteq A$, then we say that $F$ is a flippling set if the resulting graph has arc $(v,u)$ instead of arc $(u,v)$ for all $(u,v)\in F$.
In this paper we consider the following problems.
Minimum Flips For Diameter Reduction (MinFlip) Input: A diagraph $D=(V,A)$ with diameter $d_{D}$, and an integer $d\geq 2$ such that $d<d_{D}$. Goal: Find the minimum number of flips to get diameter at most $d$.
Diameter Reduction In At Most $k$ Steps ($k$-Flips) Input: A diagraph $D=(V,A)$ with diameter $d_{D}$, an integer $d\geq 2$ such that $d<d_{D}$, and a non-negative integer $k$. Goal: Decide whether it is possible to reduce the diameter to at most $d$ using at most $k$ flips.
The next claim shows the relation between the two problems.
Claim 1.
The instances of problems MinFlip and $k$-Flips are equivalent.
Proof.
We have to show that given a polynomial algorithm to either MinFlip or $k$-Flips, we can solve both of them in polynomial time. Indeed, on one hand, we can solve $k$-Flips by checking if the output of MinFlip exceeds $k$ or not. On the other hand, by solving MinFlip for $k=1,\dots,|A|$, we can solve $k$-Flips. ∎
Thus, throughout the paper we mainly concentrate on the $k$-Flips problem.
Previous Results.
In the literature the most studied problem regarding orientations is finding an oriented diameter, where an oriented diameter is the smallest diameter among all the diameters when the digraph fulfills certain properties. The input of the Oriented Diameter problem is an undirected simple graph and an integer $d\geq 0$, and the task is to find an orientation of diameter at most $d$. Chvátal and Thomassen [6] were the first ones considering the problem and they showed that finding an orientation of diameter at most 2 is $\mathsf{NP}$-hard. They also extended their result for any diameter $d\geq 4$. Further results have been considered since, but the case $d=3$ still remains open. The hardness of the problem restricted to the class of planar graphs is also not known, even though several attempts have been made (e.g. [14]).
One of the most recent improvements is by Bensmail, Duvignau and Kirgizov [3] who considered orientations with weak diameter. A digraph $D$ has $d$-weak diameter if the maximum weak distance between any pair of vertices is $d$, where a weak distance between vertices $u$ and $v$ is the minimum out of the distances from $u$ to $v$ and from $v$ to $u$ in $D$. They showed that deciding if an undirected graph has a $k$-weak orientation is $\mathsf{NP}$-complete for $k\geq 2$. They conjecture that the same holds for $d$-strong-orientations (the maximum strong distance from a vertex to another, where the strong distance between $u$ and $v$ is the maximum out of the distances from $u$ to $v$ and from $v$ to $u$ in $D$) which would complete the result of Chvátal and Thomassen for $d=3$.
Regarding Oriented Diameter in planar graphs, Eggemann and Noble [8, 9] obtained an FPT algorithm parameterized by the treewidth. Mondal, Parthiban and Rajasingh [15] studied the problem for a special type of planar graphs, called the triangular grid graphs, and they managed to give a polynomial time algorithm for this case. They also showed that the weighted version of Oriented Diameter is weakly $\mathsf{NP}$-complete even for planar graphs of bounded pathwidth, a graph parameter bounded below by treewidth. Fomin, Matamala, and Rapaport [10] considered the problem for chordal graphs, when given a bridgeless connected undirected graph $G$ and the goal is to find the smallest diameter among all the diameters of strongly connected orientations of $G$.
Ito et. al. [13] considered the problem of finding an orientation of a graph such that the digraph has maximum arc-connectivity [^2] The arc-connectivity of an undirected graph is the maximum integer $\lambda$ such that every non-empty subset $X\subsetneq V$ has at least $\lambda$ arcs leaving $X$.. Hoppenot and Szigeti [12] went further and considered a digraph $D$ that is not $k$-arc-connected but reorienting a set of arcs $F$ results on a digraph that is $k$-arc-connected. They showed that one can flip arc by arc (of $F$) without decreasing the arc-connectivity of the graph. Other problems regarding arc-connectivity and arc flipping can be found in [2].
Related results.
Finally, we want to emphasize that there is another type of flipping in the literature. Those flipping operations are related to flip graphs. A flip graph is a graph that contains combinatorial or geometric objects as vertices. An edge between two vertices indicates that those two objects can be obtained one from each other by a flip, which is an elementary operation (see e.g. [5]). For example, Aichholzer et. al. [1] considered $\alpha$-orientations of a directed graph, in which every vertex $v$ has a specific out-degree $\alpha(v)$, and a flip consists of reversing all edges of a directed cycle. Deciding whether the flip distance between two $\alpha$-orientations of flip distances on a planar graph is at most two is $\mathsf{NP}$-complete. The dual question is also $\mathsf{NP}$-hard, in which every cycle has a specified number of forward edges and a flip is the reversal of all edges contained in a minimal directed cut. Flipping an arc and the edge-flip graph of $k$-edge-connected orientations are related to orientations of undirected graphs. Ito et. al. [13] (mentioned above) also had some results regarding the edge-flip graph of $k$-arc-connected orientations whose vertices are the $k$-arc-connected orientations of the undirected graph, and two orientations are joined by an edge if and only if one is obtained from the other by a single arc flip.
Our results.
We summarize our results supplemented with the remaining (already solved or still open) cases in Table 1. Note that if $k$-Flips is solvable on planar graphs, this implies that Oriented Diameter on planar graphs is also solvable (see Section 3). On the other hand, if Oriented Diameter is $\mathsf{NP}$-hard for planar graphs, then $k$-Flips is also $\mathsf{NP}$-hard on planar graphs.
Table 1: Complexity of the problems.
Problem Parameter Fixed Complexity
$k$-Flips $\mathsf{NP}$-hard [Thm. 4]
$k$-Flips d $\mathsf{NP}$-hard [Thm. 4]
$k$-Flips k P [Thm. 12]
$k$-Flips k d $\mathsf{W[2]}$-hard [Thm. 5]
$k$-Flips on planar graphs Open
$k$-Flips on planar graphs d FPT [Thm. 15]
$k$-Flips on planar graphs k Open
$k$-Flips on cactus graphs $\mathsf{P}$ [Thm. 13]
$k$-Flips on weighted cactus graphs Weakly $\mathsf{NP}$-hard [Thm. 11]
Oriented Diameter on planar graphs Open
Oriented Diameter on planar graphs d FPT [Eggemann and Noble [8]]
We organize the remainder of the paper as follows: In Section 2 we present some basic preliminaries. In Section 3 we show that in general most related questions of this type are hard while in Section 4 we give polynomial algorithms for some special cases. The results from Section 4.3 also provide a class of counter-examples to a conjecture in the study of directed edge polytopes.
2 Preliminaries
For a given positive integer $n$, we denote by $[n]$ the set $\{1,2,\dots,n\}$.
Graph Theory.
Given an undirected graph $G=(V,E)$, a subset of vertices $S\subseteq V$ is called a dominating set if for each vertex $v$ of $G$ either $v\in S$, or $v$ has a neighbor $u$ contained in $S$, that is, $uv\in E$ and $u\in S$.
Special graph classes.
When looking for a polynomial time algorithm for an otherwise NP-hard problem, it is natural to restrict ourselves to a special family of graphs. In most cases assuming planarity makes the problem solvable. We say that a graph is planar, if we can drawn in the plane such that it does not contain any crossings. An interesting subfamily of planar graphs are the so-called cactus graphs. A cactus graph is a connected graph in which any two simple cycles have at most one common vertex. It can be represented as a tree whose nodes correspond to cycles of the cactus graph, and edges correspond to the intersection between two cycles at a vertex.
Complexity Theory.
When working on positive results for NP-hard problems, the class NP alone does not give sufficient information for intractability results. For this reason we differentiate between weakly NP-hard problems, which are hard when numbers are given in binary, and strongly NP-hard problems, which remain hard when numbers are input in unary. Similarly for FPT algorithms, intractability results usually rely on reducing to problems that are complete in some level of the $W[t]$ hierarchy [7]. For example, Clique is $W[1]$-complete and Dominating Set is $W[2]$-complete, so it is a standard assumption that neither are FPT.
Treewidth.
The treewidth of a graph measures the similarity of the graph to a tree. It was first defined by Halin but later it was rediscovered by Robertson and Seymour and also by Arnborg and Proskurowski.
For a graph $G$, a tree decomposition is a pair $(T,\mathcal{B})$ where $T$ is a tree and $\mathcal{B}$ a collection of bags $B_{v}\subseteq V(G)$, each bag corresponding to some node $v\in V(T)$. The bags have the property that for any edge of $G$, there is a bag in $\mathcal{B}$ that contains both of its endpoints. Furthermore, the bags containing $v$ form a subtree of $T$ for each vertex $v\in V(G)$. The width of $(T,\mathcal{B})$ is defined as the largest bag size of $\mathcal{B}$ minus one. The treewidth of $G$ is the minimum width $\operatorname{\mathsf{tw}}(G)$ over all tree decompositions of $G$. It is an $\mathsf{NP}$-hard graph parameter to compute, but for graphs of bounded treewidth, can be done in FPT time by the following result.
Theorem 2 (Bodlaender [4]).
For a given integer $k$ there exists an algorithm running in time $\mathcal{O}(2^{\Theta(k^{3})}n)$ that determines whether the tree-width of a graph is at most $k$, and if so, it finds a tree decomposition with tree-width at most $k$.
3 Hardness results
In this section, we revise the hardness results presented in Table 1. Let us start with our main problem, $k$-Flips, which is $\mathsf{NP}$-hard. To see this, let us consider the following problem.
Oriented Diameter Input: A simple undirected graph $G=(V,E)$, and a non-negative integer $d$. Goal: Find an orientation such that the diameter is at most $d$.
As mentioned in Section 1, this problem is known to be $\mathsf{NP}$-hard.
Theorem 3 (Chvátal – Thomassen [6]).
It is $\mathsf{NP}$-hard to decide whether an undirected graph admits an orientation of diameter 2.
Moreover, Chvátal and Thomassen extended their proof: for every $d\geq 4$, it is $\mathsf{NP}$-hard to decide whether an undirected graph admits an orientation of diameter $d$.
Theorem 4.
The problem $k$-Flips is $\mathsf{NP}$-hard.
Proof.
We show that hardness of Oriented Diameter implies hardness of $k$-Flips. It is enough to show that given a polynomial algorithm to $k$-Flips, we can solve Oriented Diameter in polynomial time. Indeed, by solving $k$-Flips with parameters $k=|A|$ and $d$, we solve Oriented Diameter with parameter $d$. ∎
Note that Oriented Diameter is not equivalent to $k$-Flips (and therefore to MinFlip) as Oriented Diameter is a special case of $k$-Flips where we fix $k=|A|$.
3.1 W[2]-hardness for number of flips
The next natural idea is to look for an FPT algorithm where the parameter is the number of flips. We show that the existence of such an algorithm is also unlikely.
Theorem 5.
Problem $k$-Flips is $\mathsf{W[2]}$-complete when parameterized by $k$.
Proof.
We make a reduction from Dominating Set which is known to be a $\mathsf{W[2]}$-complete problem (see e.g. Cygan et. al. [7]).
Dominating Set Input: An undirected graph $G=(V,E)$, and a positive integer $\ell$. Goal: Decide if there exists a dominating set of size at most $\ell$.
Let $(G,\ell)$ be an instance of Dominating Set, where $V(G)=\{v_{1},\dots,v_{n}\}$. Using $(G,\ell)$ we construct an instance of $k$-Flips as follows. For each vertex $v_{i}$, we add the vertex-gadget $H_{i}$ (see Figure 0(a)) that contains the vertices:
  • $u_{i,1},u_{i,2}$, referred as the upper vertices,
  • $d_{i,1},d_{i,2}$, referred as the down vertices,
  • $\widetilde{u}_{i,1},\widetilde{u}_{i,2}$, referred as the auxiliary upper vertices, and
  • $\widetilde{d}_{i,1},\widetilde{d}_{i,2}$, referred as the auxiliary down vertices.
We connect those vertices with the following arcs:
  • for the upper and down vertices, add arcs $(u_{i,1},u_{i,2})$, $(d_{i,1},d_{i,2})$, $(u_{i,1},d_{i,1})$, and $(d_{i,2},u_{i,2})$,
  • connect each vertex with its auxiliary pair as $(u_{i,j},\widetilde{u}_{i,j}),(\widetilde{u}_{i,j},u_{i,j})$ and $(d_{i,j},\widetilde{d}_{i,j}),(\widetilde{d}_{i,j},d_{i,j})$ for $j=1,2$,
  • connect all auxiliary vertices $x,y$ with arcs $(x,y),(y,x)$ for all pairs, except for pair $\widetilde{d}_{i,1}$ and $\widetilde{d}_{i,2}$.
Let $ij\in E(G)$, and $H_{i}$, $H_{j}$ the gadgets for vertices $i$ and $j$, respectively. We connect gadgets $H_{i}$ and $H_{j}$ in a symmetric way, meaning that the resulting arcs are the same if we change the role of $j$ and $i$ in the following description (see Figure 0(b)).
  • Add arcs $(u_{i,1},d_{j,1})$ and $(d_{j,2},u_{i,2})$, and $(u_{j,1},d_{i,1})$ and $(d_{i,2},u_{j,2})$; note that the first two arcs are symmetrical to the last two.
  • Connect all auxiliary vertices between $H_{i}$ and $H_{j}$ with arcs $(x,y),(y,x)$ for all pairs with $x\in H_{i}$ and $y\in H_{j}$; note that pair $\widetilde{d}_{i,1}$ and $\widetilde{d}_{j,2}$ can be connected.
$u_{i,1}$$u_{i,2}$$d_{i,1}$$d_{i,2}$$\widetilde{u}_{i,1}$$\widetilde{u}_{i,2}$$\widetilde{d}_{i,1}$$\widetilde{d}_{i,2}$
(a) Vertex gadget.
$u_{i,1}$$u_{i,2}$$d_{i,1}$$d_{i,2}$$u_{j,1}$$u_{j,2}$$d_{j,1}$$d_{j,2}$
(b) The connectivity of two gadgets. Auxiliary nodes are omitted for simplicity.
$u_{1,1}$$u_{1,2}$$d_{1,1}$$d_{1,2}$$\widetilde{u}_{1,1}$$\widetilde{u}_{1,2}$$\widetilde{d}_{1,1}$$\widetilde{d}_{1,2}$$u_{2,1}$$u_{2,2}$$d_{2,1}$$d_{2,2}$$\widetilde{u}_{2,1}$$\widetilde{u}_{2,2}$$\widetilde{d}_{2,1}$$\widetilde{d}_{2,2}$$u_{3,1}$$u_{3,2}$$d_{3,1}$$d_{3,2}$$\widetilde{u}_{3,1}$$\widetilde{u}_{3,2}$$\widetilde{d}_{3,1}$$\widetilde{d}_{3,2}$
(c) Example of reduction used on 5. Directed graph $H$ constructed using the complete graph $G=(V,E)$ with $V=\{v_{1},v_{2},v_{3}\}$. A solution to instance $(G,1)$ is $S=\{v_{1}\}$, while a solution to $k$-Flips is $F=\{(u_{1,1},u_{1,2})\}$ which is shown with a thick line.
Figure 1: Construction of the instance of $k$-Flips defined in 5.
We call the resulting graph $H$. An example of the construction can be seen in Figure 0(c). To conclude the reduction, set $k=\ell$. Next, we show some properties about the gadgets.
Claim 6.
$\operatorname{diam}(H_{i})=4$ holds for all $i=1,\dots,n$.
Proof of Claim 6.
First we show that the shortest path between any pair of vertices has lenght at most $4$. For auxiliary vertices the statement holds trivially. Next, we show that any auxiliary vertex $a\in\{\widetilde{u}_{i,1},\widetilde{u}_{i,2},\widetilde{d}_{i,1},\widetilde{d}_{i,2}\}$ is at distance at most 3 from any upper and down vertex $y$. First consider pair $a=\widetilde{d}_{i,1}$, $y=d_{i,2}$ and pair $a=\widetilde{d}_{i,2}$, $y=d_{i,1}$ with their paths
$(\widetilde{d}_{i,1},d_{i,1},d_{i,2}),$ (1)
$(\widetilde{d}_{i,2},\widetilde{u}_{i,1},\widetilde{d}_{i,1},d_{i,1}).$ (2)
For the rest, let $y\in\{u_{i,1},u_{i,2},d_{i,1},d_{i,2}\}$, and let $a_{y}$ be its auxiliary vertex. Then the path below has length 2 as desired
$(a,a_{y},y)$ (3)
Similarly, if we start from $x\in\{u_{i,1},u_{i,2},d_{i,1},d_{i,2}\}$ and we consider $y\in\{u_{i,1},u_{i,2},d_{i,1},d_{i,2}\}$ (except for pair $x=d_{i,2}$ and $y=d_{i,1}$), we have the path
$(x,a_{x},a_{y},y)$ (4)
which have length 3, where $a_{x}$ and $a_{y}$ are the auxiliary vertices of $x$ and $y$ respectively. Finally, for $x=d_{i,2}$ and $y=d_{i,1}$, we have the path
$(d_{i,2},u_{i,2},\widetilde{u}_{i,2},\widetilde{d}_{i,1},d_{i,1})$ (5)
with length 4. Note that this (not unique) path gives the diameter of $H_{i}$ yielding $\operatorname{diam}(H_{i})=4$. ∎
Claim 7.
$\operatorname{diam}(H)=4$.
Proof of Claim 7.
Any pair of vertices in $\{\widetilde{u}_{i,1},\widetilde{u}_{i,2},\widetilde{d}_{i,1},\widetilde{d}_{i,2},\widetilde{u}_{j,1},\widetilde{u}_{j,2},\widetilde{d}_{j,1},\widetilde{d}_{j,2}\}$ is at distance at most 2 by construction. Thus, by 6, we only need to consider the distances between vertices of different gadgets $H_{i}$ and $H_{j}$, that is, the paths from $x_{i}\in\{u_{i,1},u_{i,2},d_{i,1},d_{i,2}\}$ to $y_{j}\in\{u_{j,1},u_{j,2},d_{j,1},d_{j,2}\}$. Let $a_{i}$ and $a_{j}$ be the auxiliary vertices of $x_{i}$ and $y_{j}$, respectively. Hence, the path
$(x_{i},a_{i},a_{j},y_{j})$ (6)
has length 3, concluding the proof. ∎
A solution of the $k$-Flips instance can have a flipping set consisting on only one type of arcs.
Claim 8.
Any solution set $F^{*}$ to instance $(H,3,\ell)$ consists of a set of flipping arcs of the form $(u_{i,1},u_{i,2})$ for some indexes $i$.
Proof of Claim 8.
We check that flipping other arcs will require more than $\ell$ flips. Notice that paths of length 4 are from $d_{i,2}$ to $d_{i,1}$ (see Equation 5). Thus we can focus only on reducing the distance from $d_{i,2}$ to $d_{i,1}$.
The undirected shortest path between $d_{i,2}$ and $d_{i,1}$ has length 3 and it is given by paths:
$(d_{i,2},u_{i,2},u_{i,1},d_{i,1})$ (7)
$(d_{i,2},u_{j,2},u_{j,1},d_{i,1})\text{ for some index $j$}.$ (8)
We would like to direct at least one of those paths to make it the shortest path between those vertices. Clearly, for a connected pair $H_{i}$ and $H_{j}$, changing the direction of arcs $(d_{i,1},u_{i,1})$ and $(d_{i,2},u_{i,2})$ is not helping to get a short path from $d_{i,2}$ to $d_{i,1}$. The same holds for any arc involving auxiliary vertices, as no undirected path of length 3 contains such vertices. That is, no undirected path at distance at least 4 is of interest.
Finally, changing the direction of arc $(d_{i,1},d_{i,2})$ will only affect to gadget $H_{i}$, but not to gadget $H_{j}$.
It only remains to prove that flipping arcs $(u_{i,1},u_{i,2})$ will affect to gadget $H_{i}$ and all the other gadgets connected to it. Indeed, this follows by paths on Equation 7 and Equation 8. ∎
We can show now that the reduction is indeed correct.
Claim 9.
If $(G,\ell)$ is a Yes-instance of Dominating Set, then $(H,3,k=\ell)$ is a Yes-instance of $k$-Flips.
Proof of Claim 9.
Let $S^{*}=(s_{1},\dots,s_{\ell})$ be a dominating set of $\ell$ elements. Consider the following $\ell$ flips $F=\{(u_{i,1},u_{i,2})\,|\,v_{i}\in S^{*}\}$. Let $H^{\prime}$ be the graph after the flipping. We show that $\operatorname{diam}(H^{\prime})=3$.
First, auxiliary vertices are not modified by $F$ and the distance between any pair is still at most 3 (see paths on Equations (1), (2) and (3)). By paths of the form of Equation 4, we do not increase any path from $x\in\{u_{i,1},u_{i,2},d_{i,1},d_{i,2}\}$ to $y\in\{u_{i,1},u_{i,2},d_{i,1},d_{i,2}\}$ (except for pair $x=d_{i,2}$ and $y=d_{i,1}$) in any gadget $H_{i}$. Similarly, by Equation 6 we do not increase any path from $x_{i}\in\{u_{i,1},u_{i,2},d_{i,1},d_{i,2}\}$ to $y_{j}\in\{u_{j,1},u_{j,2},d_{j,1},d_{j,2}\}$. Also, all paths already mentioned have length 3.
We only need to verify the distance from $d_{i,2}$ to $d_{i,1}$ for all $i\in\{1,\dots,n\}$. If $i$ is an index so that $v_{i}\in S^{*}$, then the path $(d_{i,2},u_{i,1},u_{i,2},d_{i,1})$ has length 3 as desired. Otherwise, let $j$ be the index so that $v_{i}$ is adjacent to $v_{j}$ in $G$ and $v_{j}\in S^{*}$. Such index $j$ exists as $S^{*}$ is a dominating set. Thus, the path $(d_{i,2},u_{j,2},u_{j,1},d_{i,1})$ has length 3, concluding the proof. ∎
Claim 10.
If $(H,3,k=\ell)$ is a Yes-instance of $k$-Flips, then $(G,\ell)$ is a Yes-instance of Dominating Set.
Proof of Claim 10.
Let $F^{*}$ be a flipping set of $\ell$ elements, and $H^{\prime}$ the graph of diameter 3 obtained after the flipping. By 8, we know that $F^{*}$ consists of arcs $(u_{i,1},u_{i,2})$ for some gadgets $H_{i}$. Let us define $S=\{v_{i}\,|\,(u_{i,1},u_{i,2})\in F^{*}\}$. Note that $|S|=\ell$. We show that $S$ is a dominating set.
Suppose there exists an index $i$ so that $v_{i}\notin S$ and $v_{i}$ has no adjacent vertex $v_{j}\in S$. This implies that distance from $d_{i,2}$ to $d_{i,1}$ is still 4, but this contradicts the optimality of $F^{*}$. ∎
3.2 Weak $\mathsf{NP}$-hardness for Weighted Cactus Graph
In this section, we consider a special type of graphs to be named, the cactus graphs. Problem $k$-Flips for cactus graphs can be solved and we show a dynamic algorithm in Section 4.1. On the other hand, if the arcs have positive weights, the problem is weakly $\mathsf{NP}$-hard.
Theorem 11.
Problem $k$-Flips is weakly $\mathsf{NP}$-hard for weighed cactus graphs.
Proof.
We make a reduction from SubsetSum which is known to be weakly NP-hard.
SubsetSum Input: $n$ positive integers $A=\{a_{1},a_{2},\dots,a_{n}\}$ with sum $2b$. Goal: Determine if there is a subset $S\subseteq[n]$ such that $\sum_{i\in S}a_{i}=\sum_{i\not\in S}a_{i}=b$.
Given an input $(A,b)$ for SubsetSum, we construct a weighted cactus graph that is the union of $n$ cycles along a path, where each cycle contains exactly two arcs. Let $G=(V,E)$ be a graph with vertices $v_{1},v_{2},\ldots,v_{n+1}$. For all $i=1,2,\ldots,n$ we have two arcs from $v_{i}$ to $v_{i+1}$ denoted by $e_{i}$ and $f_{i}$ We put a unit weigh to arcs $e_{i}$ and weight $(a_{i}+1)$ to arcs $f_{i}$. Set $k=2n$. This is our input for $k$-Flips on a weighted cactus graph, as shown in Figure 2. Note that each arc has an orientation clockwise or counterclockwise, where originally, all arcs $e_{i}$ have a clockwise orientation, while arcs $f_{i}$ have a counterclockwise orientation (see Figure 2).
[…]
Figure 2: Instance of $k$-Flips on a weighted cactus graphs generated by an instance of SubsetSum.
If there is an orientation with diameter $(b+n)$, the distance of $v_{1}$ and $v_{n+1}$ are such that the arcs directed clockwise sum to at most $(b+n)$, and the arcs directed counterclockwise sum to at most $(b+n)$. As the total weight of the arcs is $(2b+n)$, each must be tight. If $S\subseteq[n]$ is the subset of $f$ arcs directed clockwise, we have
$\sum_{i\in S}(a_{i}+1)+\sum_{i\notin S}1$ $=\sum_{i\notin S}(a_{i}+1)+\sum_{i\in S}1$
$\sum_{i\in S}a_{i}+n$ $=\sum_{i\notin S}a_{i}+n$
$\sum_{i\in S}a_{i}$ $=\sum_{i\notin S}a_{i}=b.$
For the other direction, directing a given subset $S$ summing up $b$ clockwise and the others counterclockwise satisfies the above equations as well, making the diameter $(b+n)$. ∎
4 Algorithms
In this section we give positive results for some special cases of the problem $k$-Flips. First, we observe that simply the brute force algorithm works for any graph when $k$ is fixed.
Theorem 12.
Problem $k$-Flips is polynomially solvable if $k$ is fixed.
Proof.
Let $m$ denote the number of edges of the given graph. For all $i=1,\dots,k$ and $j=1,\dots,\binom{m}{i}$ consider all possible flipping sets $F_{ij}$. For each $F_{ij}$, using breadth first search we can calculate the diameter of the graph where the edge set $F_{ij}$ is flipped. Hence, we can decide in polynomial time whether or not the diameter can be reduced by flipping at most $k$ edges. ∎
In the remaining part of the paper we consider the problem over two special classes of graphs, namely the class of planar graphs and the class of unweighted cactus graphs.
4.1 Dynamic Algorithm for cactus graphs
When given an unweighted cactus graph, we can solve the $k$-flips problem in polynomial time.
Remark 1.
The hardness result for weighted cactus graphs does not apply here as it only showed weak NP-hardness. When given an unweighted cactus graph, we can construct a unary weighted cactus graph by contracting paths and weighting them by their lengths.
The orientations of a cactus graph are simpler to handle, as each cycle must be oriented either clockwise or counterclockwise. Furthermore, we can consider the cactus graph as a tree where nodes correspond to cycles, and edges to the intersection between these cycles.
The following dynamic programming algorithm keeps track of the minimum diameter orientation at each subtree, as well as some other intermediate data.
Theorem 13.
The $k$-Flips problem is solvable in polynomial time for unweighted cactus graphs.
Proof: Let $G_{0}=(V,A)$ be the given orientation of an underlying undirected cactus graph $G=(V,E)$ , and $k_{0},d_{0}$ the inputs to the problem. We begin by showing the problem does in fact reduce to choosing, for each cycle, whether to orient clockwise or counterclockwise.
Claim 14.
In any orientation of $G$ with a finite diameter, each of the cycles are oriented as a directed cycle.
Proof of Claim 14.
To see this, consider an orientation where a cycle $C$ does not have a cyclic orientation. Then, there is a vertex $v\in C$ with no out neighbor in $C$. If $v$ has no out neighbor outside of $C$ either, then the orientation has unbounded diameter. Thus, $v$ has some out neighbor $u\notin C$. Then, let $w\in C$ be a neighbor of $v$. By assumption, the arc $wv$ is oriented from $w$ to $v$. Now observe that there can be no path from $u$ to $w$ as the cycles are connected acyclically, so the orientation would have an infinite diameter. ∎
We can then give a flip cost $f$ for each cycle $C$ where $f(C)$ is the number of flips needed to change the orientation of $C$ in $G_{0}$ to be a clockwise orientation. The number of flips needed for a counterclockwise orientation is then $|C|-f(C)$. Let $C_{1},C_{2},\ldots,C_{k}$ be the simple cycles of $G$, and $T=(\{C_{1},C_{2},\ldots,C_{k}\},F)$ the underlying tree where an edge $\{C_{1},C_{2}\}\in F$ whenever $C_{1}$ and $C_{2}$ share a vertex. Fix an arbitrary root vertex $r$ of $T$. Given a vertex $v$ in cycles $C_{1},\ldots,C_{q}$, let $C_{v}$ denote the one furthest from $r$. In the case where this is not unique, let us use a slight abuse of notation and replace $v$ with a 0 weight cycle of length $q$ without changing the diameter. We wish to solve the following subproblem for each valid input: Given a vertex $v$, three positive integers $d_{to}$, $d_{from}$, and $k$, output the minimum diameter of an orientation of $G$ such that the maximum distance from $v$ to a node of some child of $C_{v}$ is at most $d_{to}$, the maximum distance to $v$ from a node of some child of $C_{v}$ is at most $d_{from}$, and the orientation can be reached by at most $k$ flips of the edges of $G_{0}$. We denote this minimum diameter by $D(v,d_{to},d_{from},k)$. We now describe how to compute $D(v,d_{to},d_{from},k)$ from vertices in $C_{v}$. Let $C_{v}=(v,u_{1},\ldots,u_{|C|})$ be a clockwise traversal of the vertices of $C_{v}$, and $v,u_{i_{1}},\ldots,u_{i_{\ell}}$ the vertices with degree greater than 2 in $G$. For ease of notation, we will refer to the path from $w_{1}$ to $w_{3}$ through $w_{2}$ as $(w_{1},w_{2},w_{3})$, even though it might contain more vertices. We then have to choose between orienting the edges of $C_{v}$ clockwise or counterclockwise. Let $d_{to}^{u,\text{cw}}$ ($d_{from}^{u,\text{cw}}$) be the distance from $v$ to $u$ (to $v$ from $u$) when $C_{v}$ is oriented clockwise; let $d_{to}^{u,\text{ccw}}$ ($d_{from}^{u,\text{ccw}}$) be the distance from $v$ to $u$ (to $v$ from $u$) when $C_{v}$ is oriented counterclockwise. If we orient clockwise, the to distances of the children are the following:
$d_{to}^{u_{i_{1}},\text{cw}}$ $=d_{to}-|(v,u_{1},\ldots,u_{i_{1}})|$
$d_{to}^{u_{i_{2}}\text{cw}}$ $=d_{to}-|(v,u_{1},\ldots,u_{i_{2}})|$
$\vdots$
$d_{to}^{u_{i_{\ell}}\text{cw}}$ $=d_{to}-|(v,u_{1},\ldots,u_{i_{\ell}})|.$
See Figure 3 for an example of the reasoning behind these choices. Similarly, for the maximum distances from a node of one of the children of $C_{v}$:
$d_{from}^{u_{i_{1}}\text{cw}}$ $=d_{from}-|(u_{i_{1}},u_{i_{1}+1},\ldots,v)|$
$d_{from}^{u_{i_{2}}\text{cw}}$ $=d_{from}-|(u_{i_{2}},u_{i_{2}+1},\ldots,v)|$
$\vdots$
$d_{from}^{u_{i_{\ell}}\text{cw}}$ $=d_{from}-|(u_{i_{\ell}},u_{i_{\ell}+1},\ldots,v)|.$
When orienting counterclockwise, the following parameters are needed:
$d_{to}^{u_{i_{1}},\text{ccw}}$ $=d_{to}-|(v,u_{|C|},\ldots,u_{i_{1}})|$
$d_{to}^{u_{i_{2}},\text{ccw}}$ $=d_{to}-|(v,u_{|C|},\ldots,u_{i_{2}})|$
$\vdots$
$d_{to}^{u_{i_{\ell}},\text{ccw}}$ $=d_{to}-|(v,u_{|C|},\ldots,u_{i_{\ell}})|$
$d_{from}^{u_{i_{1}},\text{ccw}}$ $=d_{from}-|(u_{i_{1}},u_{i_{1}-1},\ldots,v)|$
$d_{from}^{u_{i_{2}},\text{ccw}}$ $=d_{from}-|(u_{i_{2}},u_{i_{2}-1},\ldots,v)|$
$\vdots$
$d_{from}^{u_{i_{\ell}},\text{ccw}}$ $=d_{from}-|(u_{i_{\ell}},u_{i_{\ell}-1},\ldots,v)|.$
$u_{1}$$u_{2}$$v$
Figure 3: A picture explaining the calculations for orienting $C_{v}$. The dotted red path refers to the calculation of $d_{to}^{u_{1},\text{cw}}$ and the dashed blue $d_{to}^{u_{1},\text{ccw}}$.
With these values fixed, we can calculate the diameter of an orientation of the subgraph rooted at $v$ as the following maximum. For both orientations, we compute the maximum distance between any two nodes as
[…]
Note that $d_{to}$ and $d_{from}$ are bounded by $d_{btw}^{\text{cw}}$ and $d_{btw}^{\text{ccw}}$ as well, and that if both are finite, each value of $d_{to/from}^{u_{i},\text{cw}/\text{ccw}}$ has an orientation achieving it as well, with the appropriate number of flips. The orientation we choose is then
$D(v,d_{to},d_{from},k)=\operatorname*{arg\,min}\{d_{btw}^{\text{cw}},d_{btw}^{\text{ccw}}\}.$
Finally, we output the first case where $D(r,d,d,k)$ is less than the input $d_{0}$, as we increase $k$. If no such entry exists, then we conclude that the problem is infeasible. The actual orientation can then be found via a traceback procedure.
Correctness of this algorithm follows from the observation that we are looping through each possible consequence of each possible orientation. The running time can be bounded by the number of possible choices for $v,d_{to},d_{from},k$ times the time it takes to calculate each entry. If $n$ is the number of vertices and $m$ the number edges. There are $nm^{3}$ choices of $v,d_{to},d_{from},k$, as $k_{0}<m$. Each cycle has at most $n$ nodes , so we compute a constant times $n$ values for $d_{to/from}^{u_{i},\text{cw}/\text{ccw}}$, and take the maximum over $O(n^{2})$ terms. Thus the algorithm has complexity $O(n^{3}m^{3})$, which is polynomial in the input.
$\square$
4.2 FPT algorithm on diameter for planar graphs
We show that the Oriented Diameter problem can be solved in polynomial time on planar graphs using the following result of Eggeman and Noble.
Theorem 15 (Thm. 2.5 in [8]).
There exists an FPT algorithm parameterized by $d$, the target diameter, that inputs a planar graph $G$ and determines whether it has an orientation with diameter at most $d$.
In their paper, Eggemann and Noble used the fact that planar graphs with bounded diameter also have bounded treewidth, and then they gave an algorithm that is FPT in the treewidth of $G$. The same result holds for planar graphs of bounded oriented diameter, as the diameter of an undirected graph is bounded by the diameter of any of its orientations.
We note that this result naturally extends to the $k$-flips problem. The algorithm is very similar in nature to the one given in Section 4.1 so we can factor $k$ in similarly. For this reason, we only give a sketch of the ideas.
In the $k$-flips problem, we are given an orientation of a planar graph. We can use the underlying undirected graph, and only use the orientation when factoring in the flip distance. Let us consider a tree decomposition of this undirected planar graph with minimal treewidth, which we can construct using 2. For a given bag in the tree decomposition we keep track of all the characteristics of each vertex in the bag, where a characteristic is a summary of the distances from a given vertex to other vertices in the bag, as well as a list of vertices that are at a distance greater than $d$. We can simply add another dimension for the flip distance from the given initial orientation, and increment this appropriately to increase the complexity by a factor of $k$. Thus with only a slight modification, the algorithm of [8] solves the $k$-flips problem in time FPT in the diameter on planar graphs as well.
4.3 Directed edge polytopes
Given a symmetric graph $G=(V,A)$, its symmetric edge polytope is defined as the convex hull of the vectors $e_{v}-e_{w}\in\mathbb{R}^{V}$ where $(w,v)\in A$. In recent years, symmetric edge polytopes have become a popular object of study. It is natural to extend the definition to general directed graphs, leading to the notion of a directed edge polytope, see e.g. [11].
Studies of these objects are still limited, but observations suggested an inverse relationship between the diameter of a directed graph and the volume of the associated directed edge polytope. Cactus graphs provide a family of counter-examples to this claim.
Theorem 16.
For every integer $i\geq 8$ there exists an integer $j<i$ and a pair of directed graphs $G,H$ with diameter $i$ and $j$ respectively such that their respective directed edge polytopes have the same volume.
Proof.
We start by noticing that directed cycles correspond to standard reflexive simplices and that reversing the direction corresponds to a unimodular (and hence in particular volume preserving) transformation of the associated simplex. Further, we notice that joining two graphs at a vertex corresponds to taking the free sum of the corresponding simplices, meaning that the volumes of their corresponding directed edge polytopes are multiplied. Hence, the volume of the directed edge polytope of a directed cactus graph does not depend on the direction of the individual cycles.
Now we can construct $G$ and $H$. First assume $i$ was even. Write $i=2k$. For $G$, consider a directed path of length $k$. To every edge in that path, attach a path of length $2$ parallel to its ends and direct it the other way so that it now lies in a directed $3$-cycle. This results in a directed cactus graph of $3$-cycles with diameter $2k$. Reversing exactly one of the cycles that is attached to two other cycles results in the graph $H$ which has diameter $12k-$ 1.
If $i$ is odd, we can use an analogous construction. Write $i=2k+1$. For $G$, we start with a directed path of length $k$ again and complete every edge to a $3$-cycle, except the last one which shall be completed to a $4$-cycle instead. The construction of $H$ is analogous. ∎
While this proves that the diameter of the graph does not “control” the volume of the polytope, the question whether there exist an oriented graph with a directed edge polytope whose volume decreases as the graph undergoes a diameter reduction remains open.
5 Conclusions
The problem $k$-Flips is closely related to the problem Oriented Diameter as the latter is a special case of $k$-Flips where we set $k$ as the number of edges. We showed the complexity and gave algorithms of $k$-Flips for different settings, considering the number of flips and the target diameter as parameters. The complexity of the problem remains open on planar graphs when no parameters are given, and when $k$ is the parameter. Depending on the complexity, these two open question might help to determine the complexity for Oriented Diameter on planar graphs. That is, if there exists an algorithm for $k$-Flips on planar graphs, this algorithm will solve Oriented Diameter on planar graphs. Similarly, if Oriented Diameter is $\mathsf{NP}$-hard in planar graphs, then $k$-Flips is also $\mathsf{NP}$-hard on planar graphs.
Acknowledgements. This work was initiated at the 14th Emléktábla Workshop. We are very grateful to the organizers of the workshop. We are thankful to Lilla Tóthmérész for introducing the problem to us, and for all the fruitful discussions and comments provided. Also, we are thankful to Florian Hörsch for his valuable discussions on Section 3.1 and for bringing the reference [6] to our attention.
Grants. Panna Gehér and Lydia Mirabel Mendoza-Cadena were supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021. Lydia Mirabel Mendoza-Cadena and Daniel P. Szabo were also supported by the Hungarian National Research, Development and Innovation Office – NKFIH, grant number FK128673, and by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the ELTE TKP 2021-NKTA-62 funding scheme.